OpenStudy (anonymous):

Solve for t: t plus twenty-two point one equals twenty-three point zero eight

OpenStudy (anonymous):

t + 22.1 = 23.08

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

t + 22.1 = 23.08 - 22.1 -22.1 ______________________ t = 0.98